OpenStudy (anonymous):

solve for x. ax^n + bx^n-1 = 0

OpenStudy (anonymous):

\[ax ^{n}+bx ^{n-1}=0\]

OpenStudy (anonymous):

is it x^n(a+b^-2) = 0?

OpenStudy (anonymous):

\[x^{n-1}(ax+b)=0\] is a start

OpenStudy (anonymous):

so two answers are \[x=0, x=-\frac{b}{a}\]

OpenStudy (anonymous):

ok thank you, and I go: \[x ^{n-1} = 0 ; (ax+b) = 0\]

OpenStudy (anonymous):

i understand how to solve (ax+b) = 0, but how do i solve \[x ^{n-1}= 0\] for x